# Cubic polynomials and sums of two squares

### v1

@@ 1. Introduction
If $\mathbf{p}(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ is an irreducible cubic polynomial with $a_{2}^{2}-a_{1}$ even, and $a_{1}a_{2}-a_{0}$ odd, then
@@
Our result builds upon work of Friedlander and Iwaniec regarding values of quadratic polynomials expressible as a value of a quadratic form [2].
Our result is also related to a result of Iwaniec and Munshi [4] regarding densities of rational points on the surface $\phi(y,z)=\mathbf{p}(x)$, where $\phi$ is an irreducible quadratic form with rational coefficients, and $\mathbf{p}(x)$ is an irreducible cubic polynomial in ${\mathbb{Q}}[x]$.
@@
so that $B_{x^{3}+17}(X)\gg X^{1/2}$. Also, when $n$ is an integer we trivially have $B_{x^{3}+n^{2}}(X)\gg X^{1/2}$ as for any integer $m$, the quantity $m^{6}+n^{2}$ is a sum of two squares. In general, the author has not been able to find such polynomial substitutions for an arbitrary monic polynomial in ${\mathbb{Z}}[x]$ of degree $3$. We make the following conjecture.
Conjecture 1.2.
@@
for all $X,Y\geq 0$ and $q\in{\mathbb{N}}$. Let
$\mathcal{R}_{\theta,q}(Y_{1},Y_{2}):=\left\{(\omega_{1},\omega_{2}):\ \begin{aligned} &\omega_{i}\in{\mathbb{Z}}[\theta],\ M_{\theta}(\omega_{i})\leq Y_{1},\\ &\exists n\in{\mathbb{Z}},\ |n|\leq Y_{2},\ \omega_{1}^{2}+\omega_{2}^{2}=q^{2}(n-\theta)\end{aligned}\right\}.$
@@
Thus Theorem 1.1 follows from Lemma 1.5, Lemma 1.4 and inequality (1.1).
@@
$z-\theta=(v_{0}+v_{1}\theta+v_{2}\theta^{2})^{2}+(w_{0}+w_{1}\theta+w_{2}\theta^{2})^{2}$
@@
$\mathbf{p}(z)=\prod_{\theta:\ \mathbf{p}(\theta)=0}(z-\theta)=P_{1}P_{2}=u^{2}+v^{2}.$
@@
(3.1) $\mathcal{E}_{\theta,q}(Y_{1},Y_{2})=\left\{n:\ n\in{\mathbb{Z}},\ \#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})>0,\ |n|\leq Y_{2}\right\}.$
@@
$\mathbf{m}(\theta,q,Y_{1},Y_{2})=\max\limits_{n\in\mathbb{Z},\ \lvert n\rvert\leq Y_{2}}\#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})+1.$
@@
$\bigcup_{n\in{\mathbb{Z}},\ |n|\leq Y_{2}}\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})=$
$=\left\{(\omega_{1},\omega_{2}):\ \begin{aligned} &\omega_{i}\in{\mathbb{Z}}[\theta],\ M_{\theta}(\omega_{i})\leq Y_{1},\\ &\exists n\in{\mathbb{Z}},\ |n|\leq Y_{2},\ \omega_{1}^{2}+\omega_{2}^{2}=q^{2}(n-\theta)\end{aligned}\right\}$
@@
$\mathcal{A}_{\theta}(q^{2}(n_{1}-\theta),Y_{1})\cap\mathcal{A}_{\theta}(q^{2}(n_{2}-\theta),Y_{1})=\emptyset.$
@@
$\#\mathcal{R}_{\theta,q}(Y_{1},Y_{2})$ $=\sum_{n\in{\mathbb{Z}},\ |n|\leq Y_{2}}\#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})$
$=\sum_{n\in\mathcal{E}_{\theta,q}(Y_{1},Y_{2})}\#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})$
@@
The upper-bound follows from the triangle inequality. Let $\mathbf{p}(x)$ be the degree $3$ monic irreducible polynomial for $\theta$ with roots $\theta_{1},\theta_{2}$ and $\theta_{3}$ respectively, where without loss of generality we may assume $\theta=\theta_{1}$. Suppose that
@@
Suppose $\beta=\mu+ih$ is an algebraic integer of degree $3$ for some $h\neq 0,h\in{\mathbb{Z}}$ and $\mu\in{\mathbb{R}}$. Then it has a minimal polynomial $x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ with roots $\mu+ih$, $\mu-ih,$ and $\theta$, where $\theta$ is real. By comparing with values of coefficients we have $\theta+2\mu\in{\mathbb{Z}}$ and $2\theta\mu+\mu^{2}+h^{2}\in{\mathbb{Z}}$, therefore $\theta$ is a root of a quadratic polynomial in ${\mathbb{Q}}[x]$, but this is impossible as $\theta$ is of degree $3$ (otherwise $\beta$ will be of degree $\leq 2$).   $\sqcap\sqcup$
@@
Let $P(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$. We can consider the polynomials $P_{1}(x)$ and $P_{2}(x)$ in ${\mathbb{Z}}[i][x]$ which are defined as the binomial expansions of $P(x+i)$ and P(xi)P(x-i) respectively. It is clear that $P(x)$ is irreducible over ${\mathbb{Q}}[i]$ (otherwise ${\mathbb{Q}}[i]:{\mathbb{Q}}]\geq 3$ which is impossible), thus $P_{1}(x)$ and $P_{2}(x)$ are irreducible over ${\mathbb{Q}}[i]$. Both $P_{1}(x)$ and $P_{2}(x)$ are not polynomials in ${\mathbb{Z}}[x]$ as the $x^{2}$ coefficients of the polynomials are $3i+a_{2}$ and $(-3i+a_{2})$ respectively. Observe that the polynomial $Q(x)=P_{1}(x)P_{2}(x)$ has integer valued coefficients, thus the minimal polynomial of $\theta+i$ over the integers divides $Q(x)$.
If $Q(x)$ factors as $H_{1}(x)H_{2}(x)$, where $H_{j}(x)\in\mathbb{Q}[x]$ and $\deg(H_{j}(x))\geq 1$, then it factors similarly as $H_{1}(x)\cdot H_{2}(x)$ in ${\mathbb{Q}}[i][x]$. Since ${\mathbb{Q}}[i][x]$ is a unique factorization domain with units $c$, where $c\in{\mathbb{Q}}[i]\setminus\{0\}$, we have $H_{1}(x)=cP_{\sigma(1)}(x)$ and $H_{2}(x)=(1/c)P_{\sigma(2)}(x)$, where $\sigma:\{1,2\}\rightarrow\{1,2\}$ is a permutation map and $c\in{\mathbb{Q}}[i]$, $c\neq 0$. If $c$ is real then $H_{j}(x)\not\in{\mathbb{Z}}[x]$, which is impossible. If $c$ is complex then the $x^{3}$ coefficients of $H_{1}(x),H_{2}(x)$ are complex which is also impossible. Hence the polynomial $Q(x)$ is irreducible over $\mathbb{Z}$}.   $\sqcap\sqcup$
@@
If $\theta$ is an algebraic integer of degree $3$ then the group of units of $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ has rank of two.
@@
If $\theta$ is a real algebraic integer of degree $3$, there exists a finite set $W$ of units in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$, such that for every unit $u\in\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ there exists integers $t_{1}$ and $t_{2}$ so that
@@
Since ${\mathbb{Q}}[\theta,i]$ is closed under complex conjugation, by [1, Theorem 1] there exists a unit $q_{2}$ of $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ of infinite order with $|q_{2}|=1$ (this unit is not a torsion element). We observe that there exists integers $s_{1},s_{2},y_{1},y_{2}$ and two torsion elements $g,h\in\text{Tor}(\mathcal{O}_{{\mathbb{Q}}[\theta,i]})$ so that
@@
Observe that if $x$ is real then $q(x)$ is real, thus $q(x)\in{\mathbb{R}}[x]$. Since $p(x)$ is irreducible over ${\mathbb{Q}}[i]$, the polynomials $\prod_{j=1}^{3}\left(x-\kappa_{\theta_{1},\theta_{j}}(l_{2})\right)$ and $\prod_{j=1}^{3}\left(x-\overline{\kappa_{\theta_{1},\theta_{j}}(l_{2})}\right)$ belong to ${\mathbb{Q}}[i]$ (as both expressions can be expanded as a symmetric polynomial in $\theta_{1},\theta_{2},\theta_{3}$). Hence we deduce that $q(x)\in{\mathbb{R}}[x]\cap\left({\mathbb{Q}}[i][x]\right)$, implying that $q(x)\in{\mathbb{Q}}[x]$. By Kronecker’s theorem there exists a choice of $j\in\{2,3\}$ so that $\left|\kappa_{\theta_{1},\theta_{j}}(l_{2})\right|\neq 1$. Now note that we wish to bound from above the cardinality of the set $\mathcal{X}_{\theta}(b,l_{2},Y)$, where
@@
$\mathcal{X}_{\theta}(b,l_{2};Y)=\mathcal{X}_{\theta_{j}}(\kappa_{\theta_{1},\theta_{j}}(b),\kappa_{\theta_{1},\theta_{2}}(l_{2});Y).$
Let $\rho=|\kappa_{\theta_{1},\theta_{2}}(l_{2})|$ and $c=|\kappa_{\theta_{1},\theta_{2}}(b)|$. By lemma 3.2 there exists constants $A,B>0$ depending on $\theta$ so that
@@
By lemma 3.2 there exists positive constants $A$ and $B$ (depending only on $\theta$ and the set $W$) so that when $|w||g||l_{1}|^{t_{1}}>B\cdot Y$ or $|w||g||l_{1}|^{t_{1}}<A\cdot Y^{-2}$, the set $\mathcal{U}(g,t_{1};w,Y)$ is empty. Define the quantities
@@
We now more closely examine the set $\mathbf{S}$. From now on we will denote $\alpha$ and $\beta$ to be coprime integers such that $\alpha$ is even. Note that there exists odd integers $\mathbf{v}(\alpha,\beta)$ and $\mathbf{u}(\alpha,\beta)$ satisfying $\mathbf{u}(\alpha,\beta)\alpha-\mathbf{v}(\alpha,\beta)\beta=1$, and in particular
@@
for all integers $t$. Observe that $\mathbf{S}$ satisfies a completion property. If integer values $v_{2}=\alpha$, $u_{2}=\beta$ and $v_{1}$ are given with $v_{1}\equiv\mathbf{v}(\alpha,\beta)\pmod{2\alpha}$, then we can find a unique set of integer values $u_{0},v_{0},v_{1}$ so that
@@
for some integer $n$. Using triangle inequality we find that

### v2

@@ Abstract.
[…]
These arguments may be generalised to study the representation of irreducible monic cubic polynomials by the quadratic form $x^{2}+ny^{2}$, where $n\in\mathbb{N}$.
@@ 1. Introduction
If $\mathbf{p}(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ is an irreducible cubic polynomial which satisfies either condition 1) or condition 2), where
  • 1) $a_{2}^{2}-a_{1}$ is even, and $a_{1}a_{2}-a_{0}$ is odd
  • 2) $a_{0}\equiv 2\mod 4$, and $a_{1},a_{2}\equiv 0\mod 4$
then
@@
In a MathOverflow question, Grechuk [3] asked if there are infinitely many integers $n$ so that $n^{3}-2$ is a sum of two squares. Here we answer his question in a quantitative manner, showing that $B_{x^{3}-2}(X)\gg X^{1/3-o(1)}$. Note that non-monic cubic polynomials may never be a sum of two squares.
Our result builds upon work of Friedlander and Iwaniec regarding values of quadratic polynomials expressible as a value of a quadratic form [2]. This paper is also related to a result of Iwaniec and Munshi [4] regarding densities of rational points on the surface $\phi(y,z)=\mathbf{p}(x)$, where $\phi$ is an irreducible quadratic form with rational coefficients, and $\mathbf{p}(x)$ is an irreducible cubic polynomial in ${\mathbb{Q}}[x]$.
@@
so that $B_{x^{3}+17}(X)\gg X^{1/2}$. Also, when $n$ is an integer we trivially have $B_{x^{3}+n^{2}}(X)\gg X^{1/2}$ as for any integer $m$, the quantity $m^{6}+n^{2}$ is a sum of two squares. In general, the author has not been able to find such polynomial substitutions for an arbitrary monic polynomial in ${\mathbb{Z}}[x]$ of degree $3$. We pose the following question.
Question 1.2.
@@
[…]
$\mathcal{R}_{\theta,q}(Y_{1},Y_{2}):=\left\{(\omega_{1},\omega_{2}):\ \begin{aligned} &\omega_{i}\in{\mathbb{Z}}[\theta],\ M_{\theta}(\omega_{i})\leq Y_{1},\\ &\exists n\in{\mathbb{Z}},\ |n|\leq Y_{2},\ \omega_{1}^{2}+\omega_{2}^{2}=q(n-\theta)\end{aligned}\right\}.$
@@
Lemma 1.6.
If $\mathbf{p}(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ is an irreducible cubic polynomial with $a_{1},a_{2}\equiv 0\mod 4$, and $a_{0}\equiv 2\mod 4$, and $\theta$ is a root of $\mathbf{p}$, we have the estimate
$\#\mathcal{R}_{\theta,2}(X,X)\gg X^{1/3-o(1)}.$
Thus Theorem 1.1 follows from Lemma 1.5, Lemma 1.6, Lemma 1.4 and inequality (1.1).
@@
$q(z-\theta)=(v_{0}+v_{1}\theta+v_{2}\theta^{2})^{2}+(w_{0}+w_{1}\theta+w_{2}\theta^{2})^{2}$
@@
$q^{3}\cdot\mathbf{p}(z)=\prod_{\theta:\ \mathbf{p}(\theta)=0}q(z-\theta)=P_{1}P_{2}=u^{2}+v^{2}.$
@@
(3.1) $\mathcal{E}_{\theta,q}(Y_{1},Y_{2})=\left\{n:\ n\in{\mathbb{Z}},\ \#\mathcal{A}_{\theta}(q(n-\theta),Y_{1})>0,\ |n|\leq Y_{2}\right\}.$
@@
$\mathbf{m}(\theta,q,Y_{1},Y_{2})=\max\limits_{n\in\mathbb{Z},\ \lvert n\rvert\leq Y_{2}}\#\mathcal{A}_{\theta}(q(n-\theta),Y_{1})+1.$
@@
$\bigcup_{n\in{\mathbb{Z}},\ |n|\leq Y_{2}}\mathcal{A}_{\theta}(q(n-\theta),Y_{1})=$
$=\left\{(\omega_{1},\omega_{2}):\ \begin{aligned} &\omega_{i}\in{\mathbb{Z}}[\theta],\ M_{\theta}(\omega_{i})\leq Y_{1},\\ &\exists n\in{\mathbb{Z}},\ |n|\leq Y_{2},\ \omega_{1}^{2}+\omega_{2}^{2}=q(n-\theta)\end{aligned}\right\}$
@@
$\mathcal{A}_{\theta}(q(n_{1}-\theta),Y_{1})\cap\mathcal{A}_{\theta}(q(n_{2}-\theta),Y_{1})=\emptyset.$
@@
$\#\mathcal{R}_{\theta,q}(Y_{1},Y_{2})$ $=\sum_{n\in{\mathbb{Z}},\ |n|\leq Y_{2}}\#\mathcal{A}_{\theta}(q(n-\theta),Y_{1})$
$=\sum_{n\in\mathcal{E}_{\theta,q}(Y_{1},Y_{2})}\#\mathcal{A}_{\theta}(q(n-\theta),Y_{1})$
@@
The upper bound follows from the triangle inequality. Let $\mathbf{p}(x)$ be the degree $3$ monic irreducible polynomial for $\theta$ with roots $\theta_{1},\theta_{2}$ and $\theta_{3}$ respectively, where without loss of generality we may assume $\theta=\theta_{1}$. Suppose that
@@
Suppose $\beta=\mu+ih$ is an algebraic integer of degree $3$ for some $h\neq 0,h\in{\mathbb{Z}}$ and $\mu\in{\mathbb{R}}$. Then it has a minimal polynomial $x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ with roots $\mu+ih$, $\mu-ih,$ and $\theta$, where $\theta$ is real. By comparing with values of coefficients we have $\theta+2\mu\in{\mathbb{Z}}$ and $2\theta\mu+\mu^{2}+h^{2}\in{\mathbb{Z}}$, therefore $\theta$ is a root of a quadratic polynomial in ${\mathbb{Q}}[x]$, but this is impossible as $\theta$ is of degree $3$ (otherwise $\beta$ would be of degree $\leq 2$).   $\sqcap\sqcup$
@@
Let $P(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$. We can consider the polynomials $P_{1}(x)$ and $P_{2}(x)$ in ${\mathbb{Z}}[i][x]$ which are defined as the binomial expansions of $P(x+i)$ and P(xi)P(x-i) respectively. It is clear that $P(x)$ is irreducible over ${\mathbb{Q}}[i]$ (otherwise ${\mathbb{Q}}[i]:{\mathbb{Q}}]\geq 3$ which is impossible), thus $P_{1}(x)$ and $P_{2}(x)$ are irreducible over ${\mathbb{Q}}[i]$. Both $P_{1}(x)$ and $P_{2}(x)$ are not polynomials in ${\mathbb{Z}}[x]$ as the $x^{2}$ coefficients of the polynomials are $3i+a_{2}$ and $(-3i+a_{2})$ respectively. Observe that the polynomial $Q(x)=P_{1}(x)P_{2}(x)$ has integer-valued coefficients, thus the minimal polynomial of $\theta+i$ over the integers divides $Q(x)$.
If $Q(x)$ factors as $H_{1}(x)H_{2}(x)$, where $H_{j}(x)\in\mathbb{Q}[x]$ and $\deg(H_{j}(x))\geq 1$, then it factors similarly as $H_{1}(x)\cdot H_{2}(x)$ in ${\mathbb{Q}}[i][x]$. Since ${\mathbb{Q}}[i][x]$ is a unique factorization domain with units $c$, where $c\in{\mathbb{Q}}[i]\setminus\{0\}$, we have $H_{1}(x)=cP_{\sigma(1)}(x)$ and $H_{2}(x)=(1/c)P_{\sigma(2)}(x)$, where $\sigma:\{1,2\}\rightarrow\{1,2\}$ is a permutation map and $c\in{\mathbb{Q}}[i]$, $c\neq 0$. If $c$ is real, then $H_{j}(x)\not\in{\mathbb{Z}}[x]$, which is impossible. If $c$ is complex, then the $x^{3}$ coefficients of $H_{1}(x),H_{2}(x)$ are complex which is also impossible. Hence the polynomial $Q(x)$ is irreducible over $\mathbb{Z}$}.   $\sqcap\sqcup$
@@
If $\theta$ is an algebraic integer of degree $3$ then the group of units of $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ has rank two.
@@
If $\theta$ is a real algebraic integer of degree $3$, there exists a finite set $W$ of units in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$, such that for every unit $u\in\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ there exist integers $t_{1}$ and $t_{2}$ so that
@@
Since ${\mathbb{Q}}[\theta,i]$ is closed under complex conjugation, by [1, Theorem 1] there exists a unit $q_{2}$ of $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ of infinite order with $|q_{2}|=1$ (this unit is not a torsion element). We observe that there exist integers $s_{1},s_{2},y_{1},y_{2}$ and two torsion elements $g,h\in\text{Tor}(\mathcal{O}_{{\mathbb{Q}}[\theta,i]})$ so that
@@
Observe that if $x$ is real, then $q(x)$ is real; thus, $q(x)\in{\mathbb{R}}[x]$. Since $p(x)$ is irreducible over ${\mathbb{Q}}[i]$, the polynomials $\prod_{j=1}^{3}\left(x-\kappa_{\theta_{1},\theta_{j}}(l_{2})\right)$ and $\prod_{j=1}^{3}\left(x-\overline{\kappa_{\theta_{1},\theta_{j}}(l_{2})}\right)$ belong to ${\mathbb{Q}}[i]$ (as both expressions can be expanded as a symmetric polynomial in $\theta_{1},\theta_{2},\theta_{3}$). Hence we deduce that $q(x)\in{\mathbb{R}}[x]\cap\left({\mathbb{Q}}[i][x]\right)$, implying that $q(x)\in{\mathbb{Q}}[x]$. By Kronecker’s theorem there exists a choice of $j\in\{2,3\}$ so that $\left|\kappa_{\theta_{1},\theta_{j}}(l_{2})\right|\neq 1$. Now note that we wish to bound from above the cardinality of the set $\mathcal{X}_{\theta}(b,l_{2},Y)$, where
@@
$\mathcal{X}_{\theta}(b,l_{2};Y)=\mathcal{X}_{\theta_{j}}(\kappa_{\theta_{1},\theta_{j}}(b),\kappa_{\theta_{1},\theta_{j}}(l_{2});Y).$
Let $\rho=|\kappa_{\theta_{1},\theta_{j}}(l_{2})|$ and $c=|\kappa_{\theta_{1},\theta_{j}}(b)|$. By lemma 3.2 there exist constants $A,B>0$ depending on $\theta$ so that
@@
By Lemma 3.2 there exist positive constants $A$ and $B$ (depending only on $\theta$ and the set $W$) so that when $|w||g||l_{1}|^{t_{1}}>B\cdot Y$ or $|w||g||l_{1}|^{t_{1}}<A\cdot Y^{-2}$, the set $\mathcal{U}(g,t_{1};w,Y)$ is empty. Define the quantities
@@
We now more closely examine the set $\mathbf{S}$. From now on we will denote $\alpha$ and $\beta$ to be coprime integers such that $\alpha$ is even. Note that there exist odd integers $\mathbf{v}(\alpha,\beta)$ and $\mathbf{u}(\alpha,\beta)$ satisfying $\mathbf{u}(\alpha,\beta)\alpha-\mathbf{v}(\alpha,\beta)\beta=1$, and in particular
@@
for all integers $t$. Observe that $\mathbf{S}$ satisfies a completion property. If integer values $v_{2}=\alpha$, $u_{2}=\beta$ and $v_{1}$ are given with $v_{1}\equiv\mathbf{v}(\alpha,\beta)\pmod{2\alpha}$, then we can find a unique set of integer values $u_{0},v_{0},u_{1}$ so that
@@
for some integer $n$. Using the triangle inequality, we find that
@@
5. Proof of Lemma 1.6
We carry over useful formulas and lemmas from the previous section. We recall
$g_{0}(\mathbf{u},\mathbf{v})=$ $(u_{0}^{2}-2a_{0}u_{1}u_{2}+u_{2}^{2}a_{0}a_{2})+$
$+(v_{0}^{2}-2a_{0}v_{1}v_{2}+v_{2}^{2}a_{0}a_{2}),$
$+\left(2v_{0}v_{1}-2a_{1}v_{1}v_{2}+v_{2}^{2}(a_{1}a_{2}-a_{0})\right),$
$+(2v_{0}v_{2}+v_{1}^{2}-2a_{2}v_{1}v_{2}+v_{2}^{2}(a_{2}^{2}-a_{1})).$
$I(u_{1},u_{2},v_{1},v_{2})=u_{1}v_{2}-v_{1}u_{2}.$
We now set
$s_{1}(u_{1},u_{2},$ $v_{1},v_{2})=$
$-2+2a_{1}u_{1}u_{2}-u_{2}^{2}(a_{1}a_{2}-a_{0})+2a_{1}v_{1}v_{2}-v_{2}^{2}(a_{1}a_{2}-a_{0}),$
$s_{2}(u_{1},u_{2},$ $v_{1},v_{2})=$
$L_{0}(u_{1},u_{2},v_{1},v_{2})=v_{2}s_{1}(u_{1},u_{2},v_{1},v_{2})/2-v_{1}s_{2}(u_{1},u_{2},v_{1},v_{2})/2$
and
$M_{0}(u_{1},u_{2},v_{1},v_{2})=-u_{2}s_{1}(u_{1},u_{2},v_{1},v_{2})/2+u_{1}s_{2}(u_{1},u_{2},v_{1},v_{2})/2.$
Let $\mathbf{T}\subset{\mathbb{Z}}[\theta]\times{\mathbb{Z}}[\theta]$ contain tuples $((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2}))$ such that
  • 1) $v_{2}$ is even, $\gcd(u_{2},v_{2})=1$
  • 2) $u_{1}$ and $v_{1}$ are odd, with $I(u_{1},u_{2},v_{1},v_{2})=1$
  • 3) $u_{0}=L_{0}(u_{1},u_{2},v_{1},v_{2})$ and $v_{0}=M_{0}(u_{1},u_{2},v_{1},v_{2})$.
We verify that on $\mathbf{T}$ we have $s_{1}(u_{1},u_{2},v_{1},v_{2})\equiv 0\mod 4$, $s_{2}(u_{1},u_{2},v_{1},v_{2})\equiv 2\mod 4$. This implies that
(5.1) $u_{0}=L_{0}(u_{1},u_{2},v_{1},v_{2})\equiv 1\mod 2$
and
(5.2) $v_{0}=M_{0}(u_{1},u_{2},v_{1},v_{2})\equiv 1\mod 2.$
Lemma 5.1.
We have $\mathbf{T}\subset\mathcal{R}_{\theta,2}(\infty,\infty)$.
Proof.
By Lemma 4.1 we have
$\left((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2})\right)\in\mathcal{R}_{\theta,2}(\infty,\infty)$
if and only if all the listed conditions
  • a) $g_{0}(\mathbf{u},\mathbf{v})\equiv 0\mod 2$
  • b) $g_{1}(\mathbf{u},\mathbf{v})=-2$
  • c) $g_{2}(\mathbf{u},\mathbf{v})=0$
are satisfied. Solving the last two equations we have
$2u_{1}u_{0}+2v_{1}v_{0}=-2+2a_{1}u_{1}u_{2}-u_{2}^{2}(a_{1}a_{2}-a_{0})+2a_{1}v_{1}v_{2}-v_{2}^{2}(a_{1}a_{2}-a_{0}),$
and
where the right-hand side of the above equations are $s_{1}(u_{1},u_{2},v_{1},v_{2})$ and $s_{2}(u_{1},u_{2},v_{1},v_{2})$ respectively. This leads us to the matrix equation
$\begin{bmatrix}2u_{1}&2v_{1}\\ 2u_{2}&2v_{2}\end{bmatrix}\times\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\begin{bmatrix}s_{1}(u_{1},u_{2},v_{1},v_{2})\\ s_{2}(u_{1},u_{2},v_{1},v_{2})\end{bmatrix}$.
If $I(u_{1},u_{2},v_{1},v_{2})\neq 0$, solving the matrix equation gives us
$\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}$ $=\frac{1}{u_{1}v_{2}-v_{1}u_{2}}\begin{bmatrix}v_{2}&-v_{1}\\ -u_{2}&u_{1}\end{bmatrix}\times\begin{bmatrix}s_{1}(u_{1},u_{2},v_{1},v_{2})/2\\ s_{2}(u_{1},u_{2},v_{1},v_{2})/2\end{bmatrix}$
$=\frac{1}{u_{1}v_{2}-v_{1}u_{2}}\begin{bmatrix}v_{2}s_{1}(u_{1},u_{2},v_{1},v_{2})/2-v_{1}s_{2}(u_{1},u_{2},v_{1},v_{2})/2\\ -u_{2}s_{1}(u_{1},u_{2},v_{1},v_{2})/2+u_{1}s_{2}(u_{1},u_{2},v_{1},v_{2})/2\end{bmatrix}$
$=\frac{1}{I(u_{1},u_{2},v_{1},v_{2})}\begin{bmatrix}L_{0}(u_{1},u_{2},v_{1},v_{2})\\ M_{0}(u_{1},u_{2},v_{1},v_{2}).\end{bmatrix}$
We thus verify that elements of $\mathbf{T}$ satisfy conditions $b)$ and $c)$. Now note that on the elements of $\mathbf{T}$ we have $g_{0}(\mathbf{u},\mathbf{v})\equiv u_{0}^{2}+v_{0}^{2}\mod 2$ and by the modular equations (5.1) and (5.2) condition $a)$ is verified, and thus the proof is complete.   $\sqcap\sqcup$
The rest of the proof now follows as in the discourse after the proof of Lemma 4.2 but with $\mathbf{S}$ replaced with $\mathbf{T}$, $h_{i}$ replaced with $s_{i}$, $U_{0}$ replaced with $L_{0}$, $V_{0}$ replaced with $M_{0}$, $\mathcal{R}_{\theta,1}$ replaced with $\mathcal{R}_{\theta,2}$ and scaling absolute constants.